omunozz:

Select the correct answer from each drop-down menu. Hemoglobin Level Age Less than 25 Years 25–35 Years Above 35 Years Total Less than 9 21 32 76 129 Between 9 and 11 49 52 46 147 Above 11 69 44 40 153 Total 139 128 162 429 Based on the data in the two-way table, the probability of being 25 years or older and having a hemoglobin level above 11 is . The probability of having a hemoglobin level above 11 is . Being older than 25 years and having a hemoglobin level above 11 dependent on each other.

Bones:

The probability of being 25 years or older and having a hemoglobin level above 11 is 0.186, and the probability of having a hemoglobin level above 11 is 0.356 Probability of being 25 years or older and having a hemoglobin level above 11: To calculate this, we need to find the number of people who are 25 years or older and have a hemoglobin level above 11, which is 44 + 40 (from the "Above 35 Years" column in the "Above 11" row) = 84. Then, divide this by the total number of people in the dataset, which is 429: (84 / 429) = 0.196. Probability of having a hemoglobin level above 11: To calculate this, we need to find the total number of people with a hemoglobin level above 11, which is 69 + 44 + 40 (from the "Above 11" row) = 153. Then, divide this by the total number of people in the dataset: (153 / 429) = 0.356. Dependence: Since the probabilities of "being 25 or older and having a hemoglobin level above 11" and "having a hemoglobin level above 11" are not sign ificantly different, we can conclude that these events are not dependent on each other. Key points to remember: To calculate the probability of an event, divide the number of desired outcomes by the number of possible outcomes. Two events are considered independent if the probability of one event occurring does not affect the probability of the other event occurring.